P.T any triangle the perpendiculars from the vertices upon the opposite sides are concurrent?

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear LINTA

let the vertics of triangle are a ,b,c  and point of intersection of two perpendicular from vertics to other side is H.

so  (a-b).( H-c) =0 

 or    a.H -a.c -b.H +b.c =0 ................1

 and        (b-c).(H-a) =0

  or    b.H - b.a -c.H +c.a =0 .............2


from equation 1 and 2nd   ,   1-2

 (c-a).(H-b) =0

so points are concurrent

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin